OpenStudy (hitaro9):

Let h(x) 1/(5-x) The taylor series = sum from 0 to infinity of an(x-3)^n What is an?

OpenStudy (hitaro9):

h(x) = 1/(5-x) *

OpenStudy (anonymous):

The goal here is to write the function as one with a geometric Taylor series. Recall that \(\dfrac{1}{1-x}=\sum x^n\) for \(|x|<1\). Your desired series is centered at \(x=3\), so a slight change needs to be made to the given function: \[\frac{1}{5-x}=\frac{1}{2-(x-3)}=\frac{1}{2}\cdot\frac{1}{1-\dfrac{x-3}{2}}=\frac{1}{2}\sum_{n=1}^\infty \left(\frac{x-3}{2}\right)^n\] Simplify what you can to find \(a_n\).

OpenStudy (hitaro9):

So would it be (1/2)^(n+1)?

OpenStudy (anonymous):

Yes

OpenStudy (hitaro9):

Alright, one more small question

OpenStudy (anonymous):

Shoot

OpenStudy (hitaro9):

The series converges for |x-3| < what?

OpenStudy (hitaro9):

It's apparently not 1, so I'm a bit confused

OpenStudy (anonymous):

Applying the ratio test, we see that \[\lim_{n\to\infty}\left|\frac{(x-3)^{n+1}}{2^{n+2}}\cdot\frac{2^{n+1}}{(x-3)^n}\right|=|x-3|\lim_{n\to\infty}\left|\frac{2^{n+1}}{2^{n+2}}\right|=\frac{1}{2}|x-3|\] which converge if it's less than 1, which in turn means convergence when \(|x-3|<2\).

OpenStudy (hitaro9):

Ah. I see. Thank you so much.

OpenStudy (anonymous):

yw